40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 11662     Entry time: Sun Oct 4 13:53:30 2015
Author: jamie 
Type: Update 
Category: LSC 
Subject: SENSMAT oscillator used for EPICS tests 

I've taken over one of the SENSMAT oscillators for a test of the EPICS system.

These are the channels I've modified, with their original and current settings:

controls@donatella|~ > caget C1:LSC-OUTPUT_MTRX_7_13 C1:CAL-SENSMAT_CARM_OSC_FREQ C1:CAL-SENSMAT_CARM_OSC_CLKGAIN
C1:LSC-OUTPUT_MTRX_7_13          -1
C1:CAL-SENSMAT_CARM_OSC_FREQ    309.21
C1:CAL-SENSMAT_CARM_OSC_CLKGAIN   0
controls@donatella|~ > caget C1:LSC-OUTPUT_MTRX_7_13 C1:CAL-SENSMAT_CARM_OSC_FREQ C1:CAL-SENSMAT_CARM_OSC_CLKGAIN
C1:LSC-OUTPUT_MTRX_7_13           0
C1:CAL-SENSMAT_CARM_OSC_FREQ      0.1
C1:CAL-SENSMAT_CARM_OSC_CLKGAIN   3
controls@donatella|~ >

 

 

ELOG V3.1.3-